OpenStudy (anonymous):

(x-5/a)+1 = b solve for x and state any restrictions

OpenStudy (anonymous):

\[(x-5)/a=b-1\]\[x-5=ab-a\]\[x=ab-a+5\]

OpenStudy (lasttccasey):

Is ` (x-5/a)+1 = b ` supposed to be: \[\frac{x-5}{a}+1 = b\]or\[(x-\frac{5}{a})+1 = b\]

OpenStudy (mathmale):

@meep: lasttccasey's question is an important one. Please answer it. Then we can get this problem solved.

OpenStudy (anonymous):

It's the first one where the fraction is x-5 over a

OpenStudy (lasttccasey):

Okay then what amirreza1870 did is correct when solving for x. In this form there aren't any restrictions really but in the previous form if 'a' is zero then you'd have an undefined portion of the equation.

OpenStudy (anonymous):

Where do you get the ab-a because I don't understand that part?

OpenStudy (mathmale):

It'd be best if you were able to present your question using Equation Editor (below), because of the greater clarity that would afford you. If you can't, or don't want to, use Equation Editor, use parentheses liberally. Note how amirreza enclosed that x+5 within parentheses; you could do that yourself even without using Equation Editor: (x-5)/a+1 = b

OpenStudy (mathmale):

Starting with\[(x-5)/a=b-1\]our first step would be to eliminate the fraction on the left side. To do that, multiply both sides of the equation by 'a'. Would you do that now, please?

OpenStudy (anonymous):

I understand that part. I just don't understand where to get the -a. Because if you have ab-1 and you added 5, wouldn't that become ab+4?

OpenStudy (mathmale):

Meep: Please help me answer your question by multiplying both sides of (x-5)/a=b-1 by 'a'.

OpenStudy (anonymous):

\[x-5 = ab-1\]

OpenStudy (mathmale):

Meep, you didn't multiply the '1' by 'a'. Why?

OpenStudy (anonymous):

Ohhh Thank you for helping me find my mistake. ^^'

OpenStudy (mathmale):

Yes, this is where you were a bit off target. Can you now solve the problem?

OpenStudy (anonymous):

Yes! I got it now. Thanks so much!

OpenStudy (mathmale):

My pleasure. Bye!

OpenStudy (anonymous):

Bye!